# Talk, Listen, Connect: How Humans and AI Evaluate Empathy in Responses to Emotionally Charged Narratives

### v2

@@ 2.1.1 Human-Generated Data
Our analysis is based on data collected through an online survey conducted via Amazon MTurk in the winter of 2019 roshanaei2019empathy . This survey builds on previous IRB-approved research, where 756 videos of 126 undergraduate students were recorded. Participants were recruited using the Psychology Subject Pool and described the three best and three worst events of their lives. To ensure participant comfort with the use of their recorded videos in future studies, they have been asked to give consent for the videos in two steps. Upon arriving at the lab, participants completed a Pre-Video Recording consent form to provide consent to participate in the study. Then, after recording their videos, they filled out a Post-Video Recording consent form, explicitly indicating consent for the use of each recorded video separately. Later, each two-minute video was transcribed into text for analysis, with positive and negative labels assigned whether the story described one of the best or worst events of their life. In addition to these recorded videos, their demographic information (i.e., age, gender, race) and personality characteristics of each participant, referred to hereafter as storytellers (see Figure 1 for an example), were recorded.
@@ 2.1.4 Fine-tuning Experiment
We further explored methods of improving model performance using instruction fine-tuning to enable GPT-4o to perform more effectively. For the fine-tuning experiment, the prompt design was kept consistent with the methodology outlined above. Our data was stratified to make a balanced distribution among empathy scales, alongside persona attributes and positive and negative stories, to form a candidate training set. Finally, we make a balanced instruction training set, comprising 80 stories with similar distribution as our test set. Fine-tuning is conducted in two nested steps: first, by using the stories, that is, the stories—paired with corresponding human-annotated empathy ratings (GPT-4o FT Story-only). Second, building upon this, the model is further fine-tuned using not only the stories and empathy ratings but also additional reader-level attributes, including gender, similarity of experience with the storytellers, and self-reported empathic concern and perspective taking (GPT-4o FT All).
@@ 4 Discussion
Experimental studies in psychology have indicated several factors that play a critical role in evoking empathy. For instance, readers tend to empathize more deeply when they share similar life experiences with the storyteller krebs1975empathy ; eklund2009ve , in particular when the storyteller’s situation is particularly emotionally intense and vivid morelli2017empathy , or story touches on relatable domains such as work or home life roshanaei2019empathy . Our findings show that AI-generated responses often align with existing literature by producing higher overall empathy ratings. However, there are key factors where AI responses diverge from those of humans. AI-generated responses indicate less sensitivity to personal similarity cues, such as shared emotions, experiences, or moral values, while humans show more empathy when they feel emotionally or morally aligned with the storyteller. This highlights a key limitation of current LLMs with their lack of understanding of others’ emotions and perspectives, which is essential for authentic empathic engagement. However, our findings highlight that GPT-4o’s empathy is significantly influenced by the emotional richness, vividness, and drama of the story itself, while fails to empathize when exciting events occur. This highlights a significant gap in AI’s ability to empathize in positive circumstances, while humans tend to empathize in both positive and negative situations, including sharing joy and excitement. To further validate these findings, we examined GPT-4o’s performance separately for positive and negative stories (see Supplementary Materials S.1.3). The model demonstrated a stronger correlation with human empathy in negative contexts; however, it also tended to amplify the emotional intensity in these situations. In contrast, in positive contexts, it was less aligned with human ratings, and the degree of overestimation was smaller. This asymmetry suggests a potential bias, where LLMs may conflate emotional intensity with empathy in negative contexts, which results in inflating empathy scores for emotionally charged stories.
@@ 5 Limitations and Future Work
A second major concern in our work is the issue of personalization in empathetic AI systems. In this study, we examined the impact of incorporating persona (gender, personality, or shared experiences) and fine-tuning the models krebs1975empathy ; eklund2009ve . While personalization can improve the perceived empathy in AI systems, it also raises important ethical concerns, especially in healthcare. Healthcare data is highly sensitive, and the use of personal information to create realistic empathetic interactions requires strong safeguards to prevent misuse or breaches yalccin2019evaluating . Moreover, AI systems must be carefully designed to avoid reinforcing existing biases in healthcare. For instance, if AI is trained using biased data, it could lead to disparities in how empathy is expressed toward patients from different demographics or cultural groups, potentially worsening existing inequalities in healthcare paiva2017empathy . Given these concerns, future research must emphasize the urgent need for ethical guidelines and oversight in the development of empathetic AI, especially when personalized interactions are involved yalccin2019evaluating . In addition, the diversity of datasets should be expanded, by recruiting participants from varied cultural and demographic backgrounds to test the external validity and generalizability of findings. Furthermore, the design of empathetic AI must account for inclusivity and adaptability across diverse cultural contexts, ensuring systems can interact with a wide range of emotional expressions and cultural norms.
@@
One limitation of our study is that it relied solely on GPT-4o and a single set of prompts, potentially constraining our understanding of how different language models respond to varying inputs. Future research should explore multiple prompt variations across a broader range of models to enable a more comprehensive analysis of model behavior. Additionally, the weak alignment observed between human- and LLM-generated empathy may partly reflect the model’s limited understanding of ranking systems such as the Likert scale, despite instruction tuning. Our choice to maintain consistency in the questions posed to both humans and the AI precluded offering explicit guidance on scale interpretation, which may have contributed to the skewed empathy ratings. These findings point to valuable directions for future work, particularly in systematically examining the underlying factors and biases that shape empathy in LLMs.
@@ Abstract
S.1.3 Positive vs Negative Stories

### v3

@@ 2.1.1 Human-Generated Data
Our analysis is based on data collected through an online survey conducted via Amazon MTurk in the winter of 2019 roshanaei2024paths . This survey builds on previous IRB-approved research, where 756 videos of 126 undergraduate students were recorded. Participants were recruited using the Psychology Subject Pool and described the three best and three worst events of their lives. To ensure participant comfort with the use of their recorded videos in future studies, they have been asked to give consent for the videos in two steps. Upon arriving at the lab, participants completed a Pre-Video Recording consent form to provide consent to participate in the study. Then, after recording their videos, they filled out a Post-Video Recording consent form, explicitly indicating consent for the use of each recorded video separately. Later, each two-minute video was transcribed into text for analysis, with positive and negative labels assigned whether the story described one of the best or worst events of their life. In addition to these recorded videos, their demographic information (i.e., age, gender, race) and personality characteristics of each participant, referred to hereafter as storytellers (see Figure 1 for an example), were recorded.
@@ 2.1.4 Fine-tuning Experiment
We further explored methods of improving model performance using instruction fine-tuning to enable GPT-4o to perform more effectively. For the fine-tuning experiment, the prompt design was kept consistent with the methodology outlined above. Our data were stratified to create a balanced distribution across empathy scales, persona attributes (gender, empathic concern, perspective taking, and experience similarity), and positive and negative story tags, forming a candidate training set. Each story was treated as a single unit, with a composite categorical class constructed by concatenating the values of the stratified features. This approach ensures that stratification respects the joint distribution of these variables across training and test sets. Next, we counted the frequency of each composite class. Those with fewer than two occurrences were considered “rare” and handled separately, as they cannot be properly stratified. For non-rare classes, a stratified train-test split was applied using the composite class as the stratification criterion, and rare stories were split using a random (non-stratified) train-test split to ensure representation. Stories not assigned in either the stratified or rare splits were placed in the test set. Additionally, we checked stories in training and test sets, verified that no story appeared in both sets, and removed duplicates if necessary. Finally, we created a balanced instruction training set, comprising 80 stories with a similar distribution to our test set. Histograms of key categorical variables (e.g., tag, empathy, gender) were plotted to confirm that stratification preserved their distributions (see Supplementary materials Figure S.3). Additionally, Chi-square tests and L1-distance were used to quantify the similarity of distributions between the original and stratified datasets (see Supplementary materials Table S.1). Fine-tuning is conducted in two nested steps: first, by using the stories, that is, the stories paired with corresponding human-annotated empathy ratings (GPT-4o FT Story-only). Second, building upon this, the model is further fine-tuned using not only the stories and empathy ratings but also additional reader-level attributes, including gender, similarity of experience with the storytellers, and self-reported empathic concern and perspective taking (GPT-4o FT All).
@@ 4 Discussion
Experimental studies in psychology have indicated several factors that play a critical role in evoking empathy. For instance, readers tend to empathize more deeply when they share similar life experiences with the storyteller krebs1975empathy ; eklund2009ve , in particular when the storyteller’s situation is particularly emotionally intense and vivid morelli2017empathy , or the story touches on relatable domains such as work or home life roshanaei2024paths . Our findings show that AI-generated responses often align with existing literature by producing higher overall empathy ratings. However, there are key factors where AI responses diverge from those of humans. AI-generated responses indicate less sensitivity to personal similarity cues, such as shared emotions, experiences, or moral values, while humans show more empathy when they feel emotionally or morally aligned with the storyteller. This highlights a key limitation of current LLMs with their lack of understanding of others’ emotions and perspectives, which is essential for authentic empathic engagement.
Our findings indicate that GPT-4o’s empathy is significantly influenced by the emotional richness, vividness, and drama of the story itself, while it fails to empathize when exciting events occur. This pattern mirrors human tendencies but still indicates a larger discrepancy between the model and human empathy. To further validate these findings, we examined GPT-4o’s performance separately for positive and negative stories (see Supplementary Materials S.1.4). The model demonstrated a stronger correlation with human empathy ratings in negative contexts; however, it also tended to amplify the emotional intensity in these situations. In positive contexts, there was less alignment with human judgments, and the degree of overestimation was smaller. According to morelli2014neural ; morelli2015common there is an essential distinction between negative and positive empathy, which involves sharing differently valenced emotions (i.e., positive versus negative) and activating distinct regions in neural systems (e.g., ventromedial prefrontal cortex vs. anterior insula, dorsal anterior cingulate cortex). Negative empathy, which is often referred to as empathy in classical models, refers to resonating with another’s suffering and experiencing empathic concern and distress. In contrast, positive empathy reflects the ability to share and respond to others’ positive emotions, such as joy, pride, or success, with feelings of vicarious happiness or celebratory concern morelli2015emerging . Negative empathy typically motivates helping and supporting behaviors, whereas positive empathy supports relationship building, emotional closeness, and prosocial behavior batson1991empathic ; smith1989altruism .
Our finding shows that humans are capable of empathizing in both positive and negative contexts, like joy and excitement, as well as distress, whereas GPT-4o’s responses reveal a marked gap in its ability to empathize with positive circumstances. This asymmetry may suggest a potential bias, where LLMs may overemphasize vivid, emotionally charged distress while under-responding to positive affect. Prior research shows that people reveal greater positive empathy toward close others or in-group members, compared to negative empathy, which reflects concern and compassion, and tends to extend more broadly even toward out-group members morelli2015emerging . Compared to humans, GPT-4o’s expression of empathy appears more universal and impartial, yet it lacks the nuanced warmth and relational depth that characterize human empathy within close relationships. This pattern highlights GPT-4o’s struggle to convey the intimacy, companionship, and shared joy that underpin positive empathy in human social relationships. This limitation raises important considerations for the design and deployment of empathetic AI systems, particularly in sensitive contexts such as counseling and emotional support.
Beyond emotional resonance, our findings also underscore that cognitive empathy, the capacity to infer another person’s thoughts, intentions, and situational perspective, remains a core challenge for GPT-4o. Prior work in affective computing and AI ethics has noted that while such systems can imitate emotional expression, they struggle with perspective-taking and theory-of-mind reasoning, often simulating empathy rather than genuinely understanding it mitchell2019artificial ; bender2021stochastic ; marcus2020gpt3 . Unlike emotional empathy, which reflects affective alignment, cognitive empathy requires reasoning about others’ beliefs and intentions, abilities closely tied to the theory of mind frith2005theory . Since GPT-4o lacks embodied experience, long-term memory continuity, and grounded world models, its perspective inference remains based on linguistic associations rather than genuine comprehension. Future research could examine these alignment patterns more deeply to clarify what “empathy” entails across distinct systems of mind, human and artificial, and how each conceptualizes and interprets others’ emotions and intentions.
@@ 5 Limitations and Future Work
A second major concern in our work is the issue of personalization in empathetic AI systems. In this study, we examined the impact of incorporating persona (gender, personality, or shared experiences) and fine-tuning the models krebs1975empathy ; eklund2009ve . While personalization can improve the perceived empathy in AI systems, it also raises important ethical concerns, especially in healthcare. Healthcare data is highly sensitive, and the use of personal information to create realistic empathetic interactions requires strong safeguards to prevent misuse or breaches yalccin2019evaluating . Moreover, AI systems must be carefully designed to avoid reinforcing existing biases in healthcare. For instance, if AI is trained using biased data, it could lead to disparities in how empathy is expressed toward patients from different demographics or cultural groups, potentially worsening existing inequalities in healthcare paiva2017empathy . Given these concerns, future research must emphasize the urgent need for ethical guidelines and oversight in the development of empathetic AI, especially when personalized interactions are involved yalccin2019evaluating .
In addition, future research should expand the cultural and demographic diversity of participants to strengthen the external validity and generalizability of the findings. Because all participants in the current study were based in the United States, the observed patterns of empathy may reflect Western and individualistic norms surrounding emotional expression. Prior research has shown that various emotional experiences by individuals (e.g., anger, shame, empathy) are culturally shaped, varying across societies that emphasize collectivism versus individualism markus1991culture ; boiger2018beyond ; eichbaum2023empathy . For example, some specific aspects of self-compassion and empathy, such as Self-Kindness, Common Humanity, and Isolation, vary significantly across cultures Birkett2013SelfcompassionAE . Similarly, AI models trained mostly on Western-centric data may unintentionally encode cultural biases, which can limit their accuracy and fairness across diverse populations bender2021stochastic . In future studies, including participants from a broader range of cultural, linguistic, and socioeconomic backgrounds would help ensure that both human and AI empathy patterns are evaluated in a more globally representative context. Furthermore, the design of empathetic AI must account for inclusivity and adaptability across diverse cultural contexts, ensuring systems can interact with a wide range of emotional expressions and cultural norms.
@@
Finally, our study relies solely on GPT-4o and a single set of prompts, potentially constraining our understanding of how different language models respond to varying inputs. Future research should explore multiple prompt variations across a broader range of models to enable a more comprehensive analysis of model behavior. Additionally, the weak alignment observed between human- and LLM-generated empathy may partly reflect the model’s limited understanding of ranking systems such as the Likert scale, despite instruction tuning. Our choice to maintain consistency in the questions posed to both humans and the AI precluded offering explicit guidance on scale interpretation, which may have contributed to the skewed empathy ratings. These findings point to valuable directions for future work, particularly in systematically examining the underlying factors and biases that shape empathy in LLMs.
@@ Abstract
S.1.3 Stratification of Key Variables
Figure S.3 and Table S.1 show the distribution and statistics of the original vs. the stratified version.
Figure S.3: Distributional differences between original and stratified datasets
Variable L1 Distance Chi2 Statistic, p-value
Pos-Neg Tag 0.22 0.0000, p = 1.0000
Empathy 0.12 0.0092, p = 1.0000
Gender 0.014 0.0000, p = 1.0000
Empathic Concern 0.077 0.0035, p = 0.9983
Perspective Taking 0.039 0.0012, p = 0.9994
Experience Similarity 0.028 0.0004, p = 0.9998
Table S.1: Stratification of Key Variables for Fine-tuning.
S.1.4 Positive vs Negative Stories